# Vision-Based Mistake Analysis in Procedural Activities: A Review of Advances and Challenges

### v1

@@ VI-B Early Mistake Recognition
Early mistake recognition methods aim to identify mistakes in video segments when only an initial portion of a procedural step is observed. Unlike early action recognition [38], which seeks to infer the action being performed from a predefined action set, early mistake recognition focuses on detecting deviations from the intended action based on partial observations. This is particularly challenging as it requires attending subtle cues in the evolving action dynamics, while mistakes may manifest in diverse and often previously unseen forms.
Despite this strong relationship between the two tasks and the extensive research dedicated to the former, the latter has received comparatively less attention, with only a few works addressing it [13], [18], [21]. Qian et al. [18] target mistake detection and early mistake recognition as a video alignment problem, leveraging a Transformer-based architecture to align pairs of video sequences while jointly optimizing a video-level activity recognition loss. To facilitate early mistake recognition, they propose a baseline method based on the assumption that a reference, correct execution of an activity, $P_{\text{ref}}$, and a candidate execution, $P_{\text{cand}}$, which may contain mistakes, generally span similar time intervals. Under this assumption, the time span of both executions is partitioned into $k$ intervals and the $\ell_{2}$ distance between $P_{\text{ref}}$ and $P_{\text{cand}}$ is computed in the feature space $f$. A sudden increase in the similarity distance acts as an indicator of an unexpected event, which may correspond to a mistake in the execution process.
Wang et al., in HoloAssist [21], introduce a formulation that indirectly supports early mistake recognition via an intervention forecasting task. Framed as a temporal prediction problem, the model anticipates the need and type of intervention based on observed task progression. Interventions were delivered verbally by a domain expert monitoring the actor’s performance in real time. These serve as supervisory signals, marking critical moments where deviations from expected task execution warrant external assistance, enabling models to learn patterns indicative of upcoming failures. Finally, Peddi et al. [13], in CaptainCook4D, formulate early mistake detection in alignment with classical action anticipation paradigms [29], where a model receives a partial observation of an action segment and predicts whether it will result in a wrong or correct execution. They benchmarked this task with top-performing convolutional (SlowFast [33], X3D [39]) and transformer-based models (Omnivore [40], VideoMAE [41]) by giving them access to only the first half of an action segment. Experiments showed a significant performance drop over full-segment mistake recognition, highlighting the increased difficulty of forecasting errors under limited temporal context.
@@ VII Methodological and Supervision-Based Grouping of Mistake Analysis Approaches
Procedural activities are inherently structured, with each action contributing to a defined sequence toward a specific goal. Beyond this sequential organization, the internal characteristics of each action instance, i.e. its constituent steps, object manipulations, and evolving scene states, also determine the activity execution correctness. These intra-action dynamics often reveal subtle deviations that may not disrupt the global sequence but still signify incorrect or suboptimal execution, particularly in executional error analysis. Comprehensive mistake detection thus requires reasoning about correctness both at the macro sequence level and the micro patterns within actions. This process typically follows a multi-stage pipeline, where low-level visual perception modules (e.g., object detection [42], pose [43] and gaze estimation [44]) extract cues from video input. Mid-level modules then track interactions, model temporal dynamics, or recognize ongoing actions. Finally, high-level reasoning components compare extracted information against predefined task models or learned activity graphs to assess alignment with expected execution patterns. Deviations at the action or sequence level are flagged as potential errors.
Existing methods for mistake analysis often share core design principles, i.e. video inputs, temporal modeling, and action understanding, but differ in problem formulation, supervision use, and procedural knowledge integration. The broadest categorization follows the overarching design pipeline, shaped by the targeted error class: procedural, executional, or both. Procedural-focused methods emphasize step dependencies, action ordering [45],[37], or task progression models [18], whereas executional ones [46] focus on fine-grained motion, posture, or object manipulations, often ignoring broader context. Hybrid approaches [15] attempt to unify both perspectives, typically requiring more complex architectures capable of jointly modeling high-level task structure and low-level signals. The mistake family a method targets influences its learning objective, supervision level, and integrated priors. We group existing methods along four key pillars: (a) the specific mistake task; (b) supervision regime; (c) learning paradigm; and (d) degree of procedural structure integration, from structure-agnostic to hierarchy-aware models. Table II summarizes representative methods along these dimensions.
@@ VII-A1 Structure-Free Methods
Structure-free methods (SFM) approach procedural video understanding without imposing any prior assumptions about the underlying task structure. These models treat the input video as a continuous or disjoint sequence of frames or clips, learning to predict action labels directly. While simple and often scalable, they lack an explicit notion of step boundaries or temporal dependencies, which can limit their ability to capture procedural coherence or detect structural anomalies such as missing or disordered steps. Peddi et al. [13] evaluated several baseline methods in CaptainCook4D, focusing on structure-free approaches for supervised mistake recognition. Specifically, they examined the performance of VideoMAE [41] with a linear classifier, X3D [39], Omnivore [40] and SlowFast [33] models. These frameworks classify short video clips as correct/mistake using only visual features from fixed-length segments, without leveraging procedural context, sequential dependencies, graph-based reasoning, or alignment with predefined task templates.
Under a different setting, Lee et al. [15] utilize a two-stage, structure-free framework for procedural error detection. A temporal convolutional network first segments the input into per-frame step assignments, augmented with relational hand/object graphs to generate rich feature representations. For each step, a set of prototypes of correct executions is learned via contrastive loss, pulling frame features toward nearest prototypes and pushing them from others. At test time, each frame is scored by its cosine similarity to the closest prototype of the predicted step; low similarity indicates a procedural error (omission, addition, modification, slip, or correction).
A distinct line of research within the structure-free paradigm is proposed by Mazzamuto et al. [46], who approach mistake detection through unsupervised gaze modeling. Rather than relying on annotated action steps, temporal boundaries, or an explicit procedural structure, their method treats gaze behavior as an implicit signal of task regularity. They introduce a gaze completion task, training a model to predict future gaze distributions from past egocentric video and eye movements during correct task executions. Mistakes are detected by comparing predicted and observed gaze, with significant deviations indicating potential attentional inconsistencies in execution.
@@ VII-A2 Step-wise Models
The methods belonging to this category [21], follow a two-stage pipeline. First, a segmentation mechanism (either ground-truth or learned) yields clip-level features for each action (step). Then, each identified action step is treated as an independent unit for classification, exploiting a light-weight classification head, usually for coarse mistake detection (correct vs. incorrect). Wang et al. [21] utilize a pre-trained ViT [47] backbone to encode video clips corresponding to annotated action steps and feed the resulting per-step embeddings into a TimeSformer [48] coupled with a classification head that jointly predicts the step label and a binary “correct vs. mistake” flag. Similarly, the baseline provided in the BRIO-TA dataset [19], classifies action segments as normal or erroneous using a two-stage pipeline. First, visual features are extracted from RGB frames using spatiotemporal backbones such as I3D [32] or X3D [39]. These features are then fed into a temporal segmentation model, such as MS-TCN [35], which outputs frame-wise mistake predictions.
@@
A notable departure from these supervised paradigms is the approach in [15], which despite following a similar action localization pipeline as previous methods, it introduces a Contrastive Step Prototype Learning (CSPL) mechanism enabling step-wise mistake detection in an unsupervised manner. Instead of learning directly from both correct and erroneous executions, CSPL exploits only correct sequences. The model learns discriminative step-specific prototypes in a contrastive embedding space (via InfoNCE loss [49]), encouraging the network to produce similar embeddings for instances of the same step while pushing away different ones. At inference, each step in a test sequence is projected into this space and classified based on proximity to its corresponding prototype. Deviations from the expected prototype are flagged as errors.
@@
Figure 6: Method taxonomy based on procedural structure usage. Horizontal axis reflects the degree of structure imposed, from flat models that predict without procedural priors to structured ones that explicitly encode or reference task structure. Vertical axis captures the scope of structure, with local scope models learning short-range temporal dependencies and global scope models being capable to reason over the full procedural execution. Sketches illustrated with [8].
@@ VII-A3 Graph-Based Models
Procedural Task Graph methods (PTG): These methods utilize graphs to explicitly encode the temporal and logical structure of a procedure, where nodes typically represent discrete action steps or intermediate scene states and edges define valid transitions or causal dependencies. Such graphs may be manually annotated or automatically induced from video data [45]. They are particularly effective in modeling normal task flow and detecting anomalies, as mistake events often correspond to deviations from expected graph transitions.
@@
Seminara et al. [45], in a seminal work, investigated both approaches and their impact on the action understanding problem and the downstream task of mistake detection. For the more challenging variant of automatic graph construction, they proposed a differential learning scheme, where a neural encoder extracts frame-level features, while a probabilistic adjacency matrix, jointly optimized with step recognition, models step dependencies. To enable differentiable graph learning they proposed the Task Graph Maximum Likelihood (TGML) loss, that allows end-to-end optimization of the task graph’s adjacency matrix. It comprises two core components: (a) a positive reinforcement term, maximizing edge probabilities in the adjacency matrix for observed action transitions and (b) a contrastive penalization term, penalizing edges corresponding to action transitions absent in training sequences.
@@
Semantic Knowledge Graph methods (SKG): These methods utilize graphs to represent semantic relationships among objects, actions and their attributes within the task domain. These graphs are constructed from textual knowledge bases, external corpora, or structured ontologies. Rather than focusing on temporal ordering, semantic graphs provide contextual grounding for actions by representing affordances, object-action interactions and high-level concepts. When combined with vision models, they support reasoning about action plausibility given the current scene configuration, offering indirect support for error detection via semantic inconsistency analysis.
@@
Rule-based Graph methods (RG): These methods incorporate domain knowledge in the form of explicit procedural rules or logic, often defined by experts or extracted from formal manuals and usually are combined with the former two classes. The rules govern permissible action transitions or constraints on execution order, giving rise to graphs that represent normative task flows. Such representations are well-suited to high-precision domains such as industrial assembly or medical workflows, where deviations from the rule-defined paths are critical indicators of mistakes. Although each cluster reflects a different design philosophy for incorporating prior knowledge and modeling task regularities, a number of existing methods can belong to more than one class. An exemplar hybrid approach is the seminal work of Ding et al. [16], build a knowledge base of spatial (object-to-object relations) and temporal (orderings of spatial relations) graph structures and defines a set of logical rules in the temporal structure to detect ordering mistakes in the toy assembly and disassembly procedures in the Assembly101 [14] dataset.
@@
Template-Driven Model (TDM): TDMs incorporate external procedural templates, such as transcripts, instructional scripts, or curated step sequences, as priors to guide the understanding of complex multi-step activities. These templates define the canonical action ordering and serve as reference structures against which observed video sequences are aligned. Essentially, given a procedural template represented as an ordered action sequence and the predicted sequence, TDMs seek the optimal alignment between them under temporal consistency or monotonicity constraints, interpreting deviations as procedural or executional errors. These models leverage strong task priors encoding what should be done and in what order, enabling both action segmentation and mistake detection under weak supervision. Ghoddoosian et al. [20] employ a method based on action transcripts, using the Constrained Discriminative Forward Loss (CDFL) [50] to align frame-level predictions with expected transcripts while maximizing the decision margin between valid-invalid labels. Mistake detection is set by measuring deviations from the canonical template without error examples during training. Narasimhan et al. [51] introduce VideoTaskformer, which learns procedural structure by predicting randomly masked steps in instructional videos. Instead of aligning predictions to transcripts, it applies a masked modeling objective [52] over ordered step descriptions, using a transformer to reason over global context and identify missing or out-of-order steps disrupting procedural flow. Schoonbeek et al. [23] employ a two-stage template-driven pipeline: an Assembly State Detection (ASD) module based on YOLOv8-m [53] recognizes current assembly states from video, followed by a Procedure Step Recognition (PSR) stage mapping detected states to step sequences via non-parametric heuristics. Deviations from canonical templates, defined by task instructions, indicate procedural error presence.

### v2

@@ VI-B Early Mistake Recognition
Early mistake recognition methods aim to identify mistakes in video segments when only an initial portion of a procedural step is observed. Unlike early action recognition [kong2022human], which seeks to infer the action being performed from a predefined action set, early mistake recognition focuses on detecting deviations from the intended action based on partial observations. This is particularly challenging as it requires attending subtle cues in the evolving action dynamics, while mistakes may manifest in diverse and often previously unseen forms.
Despite this strong relationship between the two tasks and the extensive research dedicated to the former, the latter has received comparatively less attention, with only a few works addressing it [CaptainCook4D], [CSV], [HoloAssist]. Qian et al. [CSV] target mistake detection and early mistake recognition as a video alignment problem, leveraging a Transformer-based architecture to align pairs of video sequences while jointly optimizing a video-level activity recognition loss. To facilitate early mistake recognition, they propose a baseline method based on the assumption that a reference, correct execution of an activity, $P_{\text{ref}}$, and a candidate execution, $P_{\text{cand}}$, which may contain mistakes, generally span similar time intervals. Under this assumption, the time span of both executions is partitioned into $k$ intervals and the $\ell_{2}$ distance between $P_{\text{ref}}$ and $P_{\text{cand}}$ is computed in the feature space $f$. A sudden increase in the similarity distance acts as an indicator of an unexpected event, which may correspond to a mistake in the execution process.
Wang et al., in HoloAssist [HoloAssist], introduce a formulation that indirectly supports early mistake recognition via an intervention forecasting task. Framed as a temporal prediction problem, the model anticipates the need and type of intervention based on observed task progression. Interventions were delivered verbally by a domain expert monitoring the actor’s performance in real time. These serve as supervisory signals, marking critical moments where deviations from expected task execution warrant external assistance, enabling models to learn patterns indicative of upcoming failures. Finally, Peddi et al. [CaptainCook4D], in CaptainCook4D, formulate early mistake detection in alignment with classical action anticipation paradigms [lai2024human], where a model receives a partial observation of an action segment and predicts whether it will result in a wrong or correct execution. They benchmarked this task with top-performing convolutional (SlowFast [feichtenhofer2019slowfast], X3D [feichtenhofer2020x3d]) and transformer-based models (Omnivore [girdhar2022omnivore], VideoMAE [tong2022videomae]) by giving them access to only the first half of an action segment. Experiments showed a significant performance drop over full-segment mistake recognition, highlighting the increased difficulty of forecasting errors under limited temporal context.
@@ VII Methodological and Supervision-Based Grouping of Mistake Analysis Approaches
Procedural activities are inherently structured, with each action contributing to a defined sequence toward a specific goal. Beyond this sequential organization, the internal characteristics of each action instance, i.e. its constituent steps, object manipulations, and evolving scene states, also determine the activity execution correctness. These intra-action dynamics often reveal subtle deviations that may not disrupt the global sequence but still signify incorrect or suboptimal execution, particularly in executional error analysis. Comprehensive mistake detection thus requires reasoning about correctness both at the macro sequence level and the micro patterns within actions. This process typically follows a multi-stage pipeline, where low-level visual perception modules (e.g., object detection [ObjSurvey], pose [PoseSurvey] and gaze estimation [cazzato2020look]) extract cues from video input. Mid-level modules then track interactions, model temporal dynamics, or recognize ongoing actions. Finally, high-level reasoning components compare extracted information against predefined task models or learned activity graphs to assess alignment with expected execution patterns. Deviations at the action or sequence level are flagged as potential errors.
Existing methods for mistake analysis often share core design principles, i.e. video inputs, temporal modeling, and action understanding, but differ in problem formulation, supervision use, and procedural knowledge integration. The broadest categorization follows the overarching design pipeline, shaped by the targeted error class: procedural, executional, or both. Procedural-focused methods emphasize step dependencies, action ordering [seminara2024differentiable],[plini2024ti], or task progression models [CSV], whereas executional ones [mazzamuto2025gazing] focus on fine-grained motion, posture, or object manipulations, often ignoring broader context. Hybrid approaches [EgoPED] attempt to unify both perspectives, typically requiring more complex architectures capable of jointly modeling high-level task structure and low-level signals. The mistake family a method targets influences its learning objective, supervision level, and integrated priors. We group existing methods along four key pillars: (a) the specific mistake task; (b) supervision regime; (c) learning paradigm; and (d) degree of procedural structure integration, from structure-agnostic to hierarchy-aware models. Table II summarizes representative methods along these dimensions.
@@ VII-A1 Structure-Free Methods
Structure-free methods (SFM) approach procedural video understanding without imposing any prior assumptions about the underlying task structure. These models treat the input video as a continuous or disjoint sequence of frames or clips, learning to predict action labels directly. While simple and often scalable, they lack an explicit notion of step boundaries or temporal dependencies, which can limit their ability to capture procedural coherence or detect structural anomalies such as missing or disordered steps. Peddi et al. [CaptainCook4D] evaluated several baseline methods in CaptainCook4D, focusing on structure-free approaches for supervised mistake recognition. Specifically, they examined the performance of VideoMAE [tong2022videomae] with a linear classifier, X3D [feichtenhofer2020x3d], Omnivore [girdhar2022omnivore] and SlowFast [feichtenhofer2019slowfast] models. These frameworks classify short video clips as correct/mistake using only visual features from fixed-length segments, without leveraging procedural context, sequential dependencies, graph-based reasoning, or alignment with predefined task templates.
Under a different setting, Lee et al. [EgoPED] utilize a two-stage, structure-free framework for procedural error detection. A temporal convolutional network first segments the input into per-frame step assignments, augmented with relational hand/object graphs to generate rich feature representations. For each step, a set of prototypes of correct executions is learned via contrastive loss, pulling frame features toward nearest prototypes and pushing them from others. At test time, each frame is scored by its cosine similarity to the closest prototype of the predicted step; low similarity indicates a procedural error (omission, addition, modification, slip, or correction).
A distinct line of research within the structure-free paradigm is proposed by Mazzamuto et al. [mazzamuto2025gazing], who approach mistake detection through unsupervised gaze modeling. Rather than relying on annotated action steps, temporal boundaries, or an explicit procedural structure, their method treats gaze behavior as an implicit signal of task regularity. They introduce a gaze completion task, training a model to predict future gaze distributions from past egocentric video and eye movements during correct task executions. Mistakes are detected by comparing predicted and observed gaze, with significant deviations indicating potential attentional inconsistencies in execution.
@@ VII-A2 Step-wise Models
The foundational methods in this category typically follow a supervised two-stage pipeline: first segmenting the action, then classifying it as correct or incorrect. Wang et al. [HoloAssist] utilize a pre-trained ViT [dosovitskiy2020image] backbone to encode video clips corresponding to annotated action steps and feed the resulting per-step embeddings into a TimeSformer [bertasius2021space] coupled with a classification head that jointly predicts the step label and a binary “correct vs. mistake” flag. Similarly, the baseline provided in the BRIO-TA dataset [BRIOTA] utilizes spatiotemporal backbones like I3D [carreira2017quo] or X3D [feichtenhofer2020x3d] followed by temporal segmentation models (e.g., MS-TCN [farha2019ms]) to output frame-wise mistake predictions. Advancing this paradigm, Patsch et al. [patsch2025mistsense] introduce MistSense, which fuses explicit hand pose features with RGB data, each modelled via a dedicated temporal encoder and aligned via Q-Formers, to target fine-grained executional errors.
While these approaches treat the step as a monolithic unit for binary classification, recent work by Li et al. [li2025mistake] advances the STM paradigm via Mistake Attribution. Instead of a simple binary flag, their method conditions the step-wise analysis on Semantic Role Labeling [gardner2018allennlp], decomposing the textual description of an action into specific components (Predicate, Object). By cross-attending these role tokens with the video step features, the model can explicitly attribute the error to a specific semantic violation (e.g., recognizing that the object was manipulated incorrectly despite the correct action being performed). This moves modeling from coarse detection to fine-grained semantic and spatiotemporal localization.
@@
A notable departure from these supervised paradigms is the approach in [EgoPED], which despite following a similar action localization pipeline as previous methods, it introduces a Contrastive Step Prototype Learning (CSPL) mechanism enabling step-wise mistake detection in an unsupervised manner. Instead of learning directly from both correct and erroneous executions, CSPL exploits only correct sequences. The model learns discriminative step-specific prototypes in a contrastive embedding space (via InfoNCE loss [oord2018representation]), encouraging the network to produce similar embeddings for instances of the same step while pushing away different ones. At inference, each step in a test sequence is projected into this space and classified based on proximity to its corresponding prototype. Deviations from the expected prototype are flagged as errors.
Finally, a recent step-wise approach is the PECC framework [ICME], which extends the two-stage paradigm with a probabilistic treatment of step embeddings. As in prior STM models, PECC first applies a TAS backbone to obtain per-frame step labels, enforcing an explicit action segmentation. A Causal Dilated Convolution (CDC) module refines TAS features to preserve temporal dependencies and ensure causal consistency across boundaries. For each segmented step, PECC fits a Gaussian Mixture Model (GMM) over normal-execution features, yielding a probabilistic representation of step-specific appearance and motion. During inference, frames are scored by per-step log-likelihood under the corresponding GMM and flagged as errors when their likelihood deviates from the learned distribution. While it reliably identifies execution errors, it captures procedural ones only indirectly, typically when misordering or omissions cause inconsistent TAS predictions. Like all models in this category, its performance remains tied to the quality of the underlying segmentation.
@@
Figure 6: Method taxonomy based on procedural structure usage. Horizontal axis reflects the degree of structure imposed, from flat models that predict without procedural priors to structured ones that explicitly encode or reference task structure. Vertical axis captures the scope of structure, with local scope models learning short-range temporal dependencies and global scope models being capable to reason over the full procedural execution. Sketches illustrated with [openai_chatgpt].
@@ VII-A3 Graph-Based Models
Procedural Task Graph methods (PTG): These methods utilize acyclic graph structures to explicitly encode the temporal and logical structure of a procedure, where nodes typically represent discrete action steps or intermediate scene states, and, edges define valid transitions or causal dependencies. Such graphs may be manually annotated or automatically induced from video data [seminara2024differentiable]. They are particularly effective in modeling normal task flow and detecting anomalies, as mistake events often correspond to deviations from expected graph transitions.
@@
Seminara et al. [seminara2024differentiable], in a seminal work, investigated both approaches and their impact on the action understanding problem and the downstream task of mistake detection. For the more challenging variant of automatic graph construction, they proposed a differential learning scheme, where a neural encoder extracts frame-level features, while a probabilistic adjacency matrix, jointly optimized with step recognition, models step dependencies. To enable differentiable graph learning they proposed the Task Graph Maximum Likelihood (TGML) loss, that allows end-to-end optimization of the task graph’s adjacency matrix. It comprises two core components: (a) a positive reinforcement term, maximizing edge probabilities in the adjacency matrix for observed action transitions and (b) a contrastive penalization term, penalizing edges corresponding to action transitions absent in training sequences.
Building on the use of explicit procedural structures, Lee and Elhamifar [lee2025error] propose GTG2Vid, a method that formalizes the alignment of an observed video sequence to a predefined task graph as a dynamic programming problem. The method defines novel frame and node skipping dropping policies via a cumulative cost function over frame-node assignments, combining: (i) matching costs between each frame and candidate graph node (action), (ii) frame-drop costs that penalize advancing in time without transitioning to a new graph node, allowing the model to ignore ambiguous or visually uninformative frames (background actions) and (iii) node-drop costs that penalize skipping graph nodes, enabling the representation of omitted, reordered, or prematurely bypassed procedural steps. On top of this alignment, the Error Recognition Module evaluates the visual consistency of each frame with the expected embedding of its assigned graph node, identifying off-graph behavior as mistakes.
While previous methods optimize for alignment to a single path, Huang et al. [huang2025modeling] leverage the task graph to explicitly handle execution variability. In their AMNAR framework, the graph functions as a query engine that retrieves all permissible transitions given the current state, rather than enforcing a unique trajectory. This allows the model to generate adaptive ”normal” representations for multiple valid branches, effectively distinguishing between graph-compliant variations and true off-graph mistakes.
@@
Semantic Knowledge Graph methods (SKG): These methods utilize graphs to represent semantic relationships among objects, actions and their attributes within the task domain. These graphs are constructed from textual knowledge bases, external corpora, or structured ontologies. Rather than focusing on temporal ordering, semantic graphs provide contextual grounding for actions by representing affordances, object-action interactions and high-level concepts. When combined with vision models, they support reasoning about action plausibility given the current scene configuration, offering indirect support for error detection via semantic inconsistency analysis. Despite the potential of this strategy to capture richer affordance-level reasoning and to enforce semantic validity of action–object interactions, existing methods have not widely adopted it due to practical challenges: fine-grained objects in procedural tasks are difficult to detect and track reliably, generic detectors rarely cover task-specific components, and grounding semantic relations in visual representations requires substantial task-specific annotation and model fine-tuning. These limitations make robust semantic reasoning difficult to achieve in real-world procedural video settings.
@@
Rule-based Graph methods (RG): These methods incorporate domain knowledge in the form of explicit procedural rules or logic, often defined by experts or extracted from formal manuals and usually are combined with the former two classes. The rules govern permissible action transitions or constraints on execution order, giving rise to graphs that represent normative task flows. Such representations are well-suited to high-precision domains such as industrial assembly or medical workflows, where deviations from the rule-defined paths are critical indicators of mistakes. Although each cluster reflects a different design philosophy for incorporating prior knowledge and modeling task regularities, a number of existing methods can belong to more than one class. An exemplar hybrid approach is the seminal work of Ding et al. [ding2023mistakecountsassembly], build a knowledge base of spatial (object-to-object relations) and temporal (orderings of spatial relations) graph structures and defines a set of logical rules in the temporal structure to detect ordering mistakes in the toy assembly and disassembly procedures in the Assembly101 [Assembly101] dataset.
@@
Template-Driven Model (TDM): TDMs incorporate external procedural templates, such as transcripts, instructional scripts, or curated step sequences, as priors to guide the understanding of complex multi-step activities. These templates define the canonical action ordering and serve as reference structures against which observed video sequences are aligned. Essentially, given a procedural template represented as an ordered action sequence and the predicted sequence, TDMs seek the optimal alignment between them under temporal consistency or monotonicity constraints, interpreting deviations as procedural or executional errors. These models leverage strong task priors encoding what should be done and in what order, enabling both action segmentation and mistake detection under weak supervision. Ghoddoosian et al. [ATA] employ a method based on action transcripts, using the Constrained Discriminative Forward Loss (CDFL) [li2019weakly] to align frame-level predictions with expected transcripts while maximizing the decision margin between valid-invalid labels. Mistake detection is set by measuring deviations from the canonical template without error examples during training. Narasimhan et al. [narasimhan2023learning] introduce VideoTaskformer, which learns procedural structure by predicting randomly masked steps in instructional videos. Instead of aligning predictions to transcripts, it applies a masked modeling objective [devlin2019bert] over ordered step descriptions, using a transformer to reason over global context and identify missing or out-of-order steps disrupting procedural flow. Schoonbeek et al. [IndustReal] employ a two-stage template-driven pipeline: an Assembly State Detection (ASD) module based on YOLOv8-m [yolov8_ultralytics] recognizes current assembly states from video, followed by a Procedure Step Recognition (PSR) stage mapping detected states to step sequences via non-parametric heuristics. Deviations from canonical templates, defined by task instructions, indicate procedural error presence.